Session: Feb 2001 (session 1)
Wavelength: 6.0 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Pisanu at Mon Feb 12 03:58:09 MET 2001) Minor failures
From 01:30 UT, problems with the antenna.

Effelsberg (A. Kraus at Mon Feb 12 09:50:38 MET 2001) Minor failures
The scan at UT0050-0055 was lost - problems with the telescope control computer.

Torun (K.Borkowski at Mon Feb 12 09:53:25 MET 2001) Success
No known problems.

Onsala (John Conway at Mon Feb 12 10:08:33 MET 2001) Success
No known problems

Jodrell_Bank (Alastair Gunn at Mon Feb 12 13:22:39 MET 2001) Minor failures
FS restarted at approximately 19:00 UT. No loss of data. No other known problems.

Westerbork (Tony Foley at Mon Feb 12 15:17:31 MET 2001) Minor failures
2 cases where trying to observe a source just outside hour angle range (about 18:39 and 21:56) triggered a TMS bug. We lost 18:38 to 19:43 and 21:56 to 22:36.

Noto (C.Trigilio at Thu Feb 15 18:42:46 MET 2001) Minor failures
Formatter delay of 1 sec found and corrected at 20:01 UT.